Suppose f(x)={kx + 7 ; x>= 2 ,x^2+19 ; x < 2' , For what value of k is lim x→2 f(x) defined?


1
Expert's answer
2021-12-02T07:23:23-0500

"f(x) = \begin{cases}
 x^2+19 &x<2 \\
 kx + 7 &x\geq 2
\end{cases}"

"\lim\limits_{x\to 2^-}f(x)=\lim\limits_{x\to 2^-}(x^2+19)=(2)^2+19=23"

"\lim\limits_{x\to 2^+}f(x)=\lim\limits_{x\to 2^+}(kx+7)=2k+7"

If "\lim\limits_{x\to 2}f(x)" is defined, then


"\lim\limits_{x\to 2^-}f(x)=\lim\limits_{x\to 2^+}f(x)=\lim\limits_{x\to 2}f(x)"

Hence


"23=2k+7=>k=8"



"\lim\limits_{x\to 2}f(x)" is defined, if "k=8."